The perimeter of a rectangular fence must be at least 400 feet and no more than 480 feet. If the length of the fence must be three times the width, what is the range of values allowed for the width of the fence?
Answer by addingup(3677)

Perimeter of the rectangle: 2Length + 2Width
In this problem:
L = 3W
In the formula for the perimeter:
2L + 2W
substitute for L
2(3W) + 2W = 400
6W + 2W = 400
8W = 400
W = 50
--------------------
Now you do
2(3W)+2W = 480
and the result you get gives you the max width, your range is between 50 and the answer you get.